OpenStudy (anonymous):

Write a rational expression with excluded values -2 and 2. Explain how you found your answer. I will reward medals!~

OpenStudy (anonymous):

(x-2)(x+2)=x^2+2x-2x-4=x^2-4 so x^2-4

OpenStudy (anonymous):

So x^2-4 is your answer?

OpenStudy (anonymous):

more like something/x^2-4 example (x^2)/(x^2-4)

OpenStudy (anonymous):

So why it would be that?